OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t − 6(1 + t) −t − 6 −t + 6 t + 6 6t − 6

OpenStudy (king):

-6(1+t)=-6-6t 5t-6-6t =-6-t -t-6